OpenStudy (anonymous):

Solve. –18 = 6t (Points : 1) 12 24 –3 04_03_LQ_5d.gif

OpenStudy (anonymous):

t = -3

OpenStudy (anonymous):

you have to divide both sides by 6

OpenStudy (academicgurusinc):

Solve. –18 = 6t (Points : 1) 12 24 –3 Step 1: Identify the goal. The goal is to solve for t. Step 2: What needs to be done in order to isolate t is to get rid of he 6, which is multiplying t. You get rid of the 6 by dividing it from both sides of the equation: Original question : –18 = 6t Divide both sides of equation by 6: (1/6)(-18) = (1/6)(6)(t) Result:-3 =t If you need additional clarification, feel free to drop us a line at info@academicgurusinc.com. Follow us on twitter (@Academic_Gurus) and Facebook (AcademicGurusInc). Leave us a comment and let us know if we were able to help. Best Regards, AcademicGurusInc

OpenStudy (anonymous):

super overboard for a easy question ^